OpenStudy (anonymous):

cos(50) = t; determine cos(25) in terms of t

hartnn (hartnn):

Hi , Welcome to OpenStudy :) Do you know the standard formula for cos 2x ? in terms of cos x

OpenStudy (anonymous):

thank you. I know the double angle formula for cos2x

hartnn (hartnn):

good, so what is it ?

OpenStudy (anonymous):

cos^2x -sin^2x

hartnn (hartnn):

do you know any other form , which has only cos term ?

OpenStudy (anonymous):

i know 2cos^2x-1

hartnn (hartnn):

yes, thats what i was asking :) so, \(\cos 2x = 2\cos^2x-1\) can you isolate cos x from that ?

OpenStudy (anonymous):

oh sorry :) wouldnt cosx = square root( (1+ cos2x)/2) ?

hartnn (hartnn):

correct ! good :) so, in cosx = square root( (1+ cos2x)/2 just put x= 25 what u get ?

OpenStudy (anonymous):

oh my goodness, i was so silly not to see that! you get square root((1+t)/2). Thank you so much for the help :) i really appreciate it!

hartnn (hartnn):

your welcome ^_^